Quantitative Problem: Barton Industries expects next year's annual dividend, D1, to be $2.40 and it expects dividends to grow at a constant rate gL = 4.8%. The firm's current common stock price, P0, is $24.00. If it needs to issue new common stock, the firm will encounter a 5.6% flotation cost, F. Assume that the cost of equity calculated without the flotation adjustment is 14.8% and the cost of old common equity is 14.2%. What is the flotation cost adjustment that must be added to its cost of retained earnings? Do not round intermediate calculations. Round your answer to two decimal places.

  %

What is the cost of new common equity considering the estimate made from the three estimation methodologies? Do not round intermediate calculations. Round your answer to two decimal places.

  %
Answer #1

Details given

D1 = $ 2.40

Dg = 4.8%

P0 = 24

F = 5.6%

Cost of equity without flotation costs = 14.8%

Cost of Old common equity = 14.2%

1) Cost of new common equity

Cost of equity with flotation costs

= D1 / P0(1-F)+Dg

= 2.4 / 24(1-5.6%)+4.8%

= 2.4 / 24 (0.944)+4.8%

= 2.4 / 22.66+4.8%

=10.59%+4.8%

=15.39%

2) Flotation cost adjustment to retained earnings

Flotation adjustment is the difference between cost new equity including Flotation costs less Cost of equity without flotation costs

= 15.39%-14.8%

= 0.59%